# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

### v1

@@ 4.3 Automatic Covariate Balancing
4.4 Constrained Optimization Problem
where we replace $\alpha_{0}$ with $\widehat{\alpha}$. Then, it holds that
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\partial g\big(\alpha(X_{i})\big)-m(W_{i},(\partial g)\circ\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,k.$
4.5 Linearity for the Basis Functions
Next, we consider the case where
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{k}_{j=1}{\bm{\beta}}_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right).$
We consider the following constrained optimization problem:
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)-m(W_{i},(\partial g)\circ\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,k.$
4.6 Dual Formulation
Using Lagrange multipliers ${\bm{\beta}}_{j}\in{\mathbb{R}}$ ($j=1,2,\dots,k$), the constrained optimization problem can be written as
$\min_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}\sup_{{\bm{\beta}}\in{\mathbb{R}}^{k}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})+\sum^{k}_{j=1}{\bm{\beta}}_{j}\left(\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-\operatorname{sign}({\bm{\beta}}_{j})\lambda\right)\right\}.$
The dual problem of the above constrained problem is
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{k}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})+\sum^{k}_{j=1}\left({\bm{\beta}}_{j}\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-|{\bm{\beta}}_{j}|\lambda\right)\right\}.$
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{k}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}\left(g(\alpha_{i})+\sum^{k}_{j=1}\left({\bm{\beta}}_{j}\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-|{\bm{\beta}}_{j}|\lambda\right)\right).$
From $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{k}_{j=1}{\bm{\beta}}_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)$, we have
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{k}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{i})-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda|{\bm{\beta}}|_{1}\right\}.$
Consider the problem
$\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{i})-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda|{\bm{\beta}}|_{1}\right\}.$
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{k}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{{\bm{\beta}}}(X_{i}))-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)-\lambda|{\bm{\beta}}|_{1}\right\}.$
This is equivalent to
$\min_{{\bm{\beta}}\in{\mathbb{R}}^{k}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha_{i})+\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)-m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda|{\bm{\beta}}|_{1}\right\}.$
@@ 4.7 Choice of Loss and Link Functions
For the covariate balancing property, the linearity assumption $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{k}_{j=1}\beta_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)$ is crucial. This assumption depends on the choice of the function $g$ in the Bregman divergence, which corresponds to the loss function, and on the choice of the link function.
For example, for SQ-Riesz regression, UKL-Riesz regression, and BP-Riesz regression, the derivative of $g$ is given as
  • Squared loss (SQ-Riesz): $\partial g^{\text{SQ}}(\alpha)=\alpha-1$.
@@
  • BP divergence loss (BP-Riesz): $\partial g^{\text{BP}}(\alpha)=\operatorname{sign}(\alpha)\Big(1+\frac{1}{\gamma}\Big)\Big(\big(|\alpha|-C\big)^{\gamma}-1\Big),\qquad\mathrm{dom}\left(\partial g^{\text{BP}}\right)=\{\alpha\in{\mathbb{R}}:|\alpha|\geq C\}$ for $\gamma>0$, $\alpha<-C$, and $\alpha>C$. We introduce the following branchwise maps:
    $\partial g^{\text{BP}}_{+}(\alpha)\coloneqq\Big(1+\frac{1}{\gamma}\Big)\Big((\alpha-C)^{\gamma}-1\Big),\qquad\alpha\in[C,\infty),$
    $\partial g^{\text{BP}}_{-}(\alpha)\coloneqq-\Big(1+\frac{1}{\gamma}\Big)\Big((-\alpha-C)^{\gamma}-1\Big),\qquad\alpha\in(-\infty,-C],$
@@
  • Squared loss (SQ-Riesz):
    $\left(\partial g^{\text{SQ}}\right)^{-1}(v)\coloneqq v+1\qquad(v\in{\mathbb{R}}).$
@@
    $\left(\partial g^{\text{BP}}_{+}\right)^{-1}(v)$ $=C+\Big(1+\frac{v}{k}\Big)^{1/\gamma},\qquad(v\geq-k),$
    $\left(\partial g^{\text{BP}}_{-}\right)^{-1}(v)$ $=-C-\Big(1-\frac{v}{k}\Big)^{1/\gamma},\qquad(v\leq k),$
    for $k\coloneqq 1+1/\gamma$.
These inverse functions suggest the following link functions, under which we can attain the linearity property:
  • Squared loss (SQ-Riesz): We use the identity link function,
    $\alpha_{{\bm{\beta}}}(X)=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)={\bm{\phi}}(X)^{\top}{\bm{\beta}}+1.$
    $\alpha_{{\bm{\beta}}}(X)=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)=\xi(X)\Big(C+\exp\big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big)-(1-\xi(X))\Big(C+\exp\big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big).$
  • BP divergence loss (BP-Riesz):
    $\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)=\xi(X)\left(C+\Big(1+\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\gamma}\right)-(1-\xi(X))\left(C+\Big(1-\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\gamma}\right),$
    for $k\coloneqq 1+1/\gamma$.
@@ 4.8 SQ-Riesz regression with a Linear Link Function
For this model, using the squared loss (SQ-Riesz regression, or standard Riesz regression) yields automatic covariate balancing. Specifically, under a linear model, the dual formulation implies that SQ-Riesz regression is equivalent to solving
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}(|\alpha_{i}|-C)\log\big(|\alpha_{i}|-C\big)-(|\alpha_{i}|+C)\log(|\alpha_{i}|+C)$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,k.$
@@ 4.9 UKL-Riesz Regression with a Log Link Function
For this model, using the UKL divergence (UKL-Riesz regression) yields automatic covariate balancing. Specifically, under this specification, the dual formulation implies that UKL-Riesz regression is equivalent to solving
@@
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\Big(\xi(X_{i})\phi(X_{i})-(1-\xi(X_{i}))\phi(X_{i})\Big)-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)\right|\leq\lambda\qquad j=1,\dots,k.$
@@
Special case of the log link function
@@
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\phi(X_{i})-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)\right|\leq\lambda\qquad\text{for}\ \ j=1,\dots,k.$
@@
4.10 BP-Riesz Regression and a Power Series Function
@@
Let $\gamma\in(0,\infty)$ and define $k\coloneqq 1+1/\gamma$. Consider the following model for the Riesz representer:
$\alpha_{\bm{\beta}}(X)=\xi(X)\Big(C+\Big(1+\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\gamma}\Big)-(1-\xi(X))\Big(C+\Big(1-\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\gamma}\Big),$ (3)
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function and $\xi\colon{\mathcal{X}}\to\{0,1\}$ selects the branch. We call (3) a power series link function because, on each branch and for $|{\bm{\phi}}(X)^{\top}{\bm{\beta}}|<k$, the map
$t\mapsto\Big(1\pm\frac{t}{k}\Big)^{1/\gamma}$
admits a power series expansion around $t=0$. In particular, near $t=0$ we have
$\Big(1+\frac{t}{k}\Big)^{1/\gamma}=1+\frac{t}{\gamma k}+O(t^{2})=1+\frac{t}{1+\gamma}+O(t^{2}),$
so the link behaves approximately linearly around the origin, while still allowing nonlinear growth away from the origin.
@@
Under (3), the dual characterization implies that BP-Riesz regression returns the minimum BP-loss solution among approximately balancing models. In particular, BP-Riesz regression is equivalent to solving a constrained problem of the form
@@
The power series link (3) provides a continuous bridge between the specifications in the previous subsections. As $\gamma\to 0$, we have $k=1+1/\gamma\to\infty$ and
$\Big(1+\frac{t}{k}\Big)^{1/\gamma}=\Big(1+\gamma t+o(\gamma)\Big)^{1/\gamma}\to\exp(t),$
@@
so (3) reduces to the log link form used for UKL-Riesz regression. At $\gamma=1$, the BP loss reduces to the squared loss, and the link becomes an affine transformation of the linear index around the origin, which connects BP-Riesz regression to SQ-Riesz regression up to reparameterization.
@@
This interpolation perspective is also consistent with the robustness interpretation of the BP divergence (Basu et al., 1998; Sugiyama et al., 2012). Smaller $\gamma$ makes the objective closer to a KL-type criterion, which can be efficient under correct specification, while larger $\gamma$ yields behavior closer to squared loss and is typically more robust to misspecification and extreme weights.

### v2

@@ 4.3 Choice of Loss and Link Functions
For the automatic covariate balancing property, the linearity assumption
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$
or $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)$ are crucial. These assumptions depend on the choice of the function $g$ in the Bregman divergence, which corresponds to the loss function, and on the choice of the link function.
For example, for SQ-Riesz regression, UKL-Riesz regression, and BP-Riesz regression, the following choice of Riesz representer models (link functions) guarantee the linearity assumption.
  • Squared distance (SQ-Riesz): We use the identity link function,
    $\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$
    $=({\bm{\phi}}(X)^{\top}{\bm{\beta}}+C)/2.$
    $\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$
    $=\xi(X)\Big(C+\exp\big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big)-(1-\xi(X))\Big(C+\exp\big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big).$
  • BP divergence loss (BP-Riesz):
    $\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$
    $=\xi(X)\left(C+\Big(1+\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\delta}\right)-(1-\xi(X))\left(C+\Big(1-\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\delta}\right),$
    for $k\coloneqq 1+1/\delta$.
Justification.
We justify the above choices of the link functions by showing that under them, the automatic covariate balancing property holds. The derivative of $g$ is given as
  • Squared distance (SQ-Riesz): $\partial g^{\text{SQ}}(\alpha)=2(\alpha-C)$.
@@
  • BP divergence loss (BP-Riesz): $\partial g^{\text{BP}}(\alpha)=\operatorname{sign}(\alpha)\Big(1+\frac{1}{\delta}\Big)\Big(\big(|\alpha|-C\big)^{\delta}-1\Big),\qquad\mathrm{dom}\left(\partial g^{\text{BP}}\right)=\{\alpha\in{\mathbb{R}}:|\alpha|\geq C\}$ for $\delta>0$, $\alpha<-C$, and $\alpha>C$. We introduce the following branchwise maps:
    $\partial g^{\text{BP}}_{+}(\alpha)\coloneqq\Big(1+\frac{1}{\delta}\Big)\Big((\alpha-C)^{\delta}-1\Big),\qquad\alpha\in[C,\infty),$
    $\partial g^{\text{BP}}_{-}(\alpha)\coloneqq-\Big(1+\frac{1}{\delta}\Big)\Big((-\alpha-C)^{\delta}-1\Big),\qquad\alpha\in(-\infty,-C],$
@@
  • Squared distance (SQ-Riesz):
    $\left(\partial g^{\text{SQ}}\right)^{-1}(v)\coloneqq(v+C)/2\qquad(v\in{\mathbb{R}}).$
@@
    $\left(\partial g^{\text{BP}}_{+}\right)^{-1}(v)$ $=C+\Big(1+\frac{v}{k}\Big)^{1/\delta},\qquad(v\geq-k),$
    $\left(\partial g^{\text{BP}}_{-}\right)^{-1}(v)$ $=-C-\Big(1-\frac{v}{k}\Big)^{1/\delta},\qquad(v\leq k),$
    for $k\coloneqq 1+1/\delta$.
These inverse functions suggest the above link functions.
@@ 4.4 SQ-Riesz regression with a Linear Link Function
For this model, using the squared distance (SQ-Riesz regression, or standard Riesz regression) yields automatic covariate balancing. Specifically, under a linear model, if we use $\ell_{1}$-penalty in SQ-Riesz regression, the dual formulation implies that SQ-Riesz regression is equivalent to solving
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}(\alpha_{i}-C)^{2}$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,p.$
@@ 4.5 UKL-Riesz Regression with a Log Link Function
For this model, using the UKL divergence (UKL-Riesz regression) yields automatic covariate balancing. Specifically, under this specification, if we use $\ell_{1}$-penalty in UKL-Riesz regression, the dual formulation implies that UKL-Riesz regression is equivalent to solving
@@
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\Big(\xi(X_{i})\phi(X_{i})-(1-\xi(X_{i}))\phi(X_{i})\Big)-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)\right|\leq\lambda\qquad j=1,\dots,p.$
@@
Special case of the log link function.
@@
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\phi_{j}(X_{i})-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)\right|\leq\lambda\qquad\text{for}\ \ j=1,\dots,p.$
@@
4.6 BP-Riesz Regression and a Power Link Function
@@
Let $\delta\in(0,\infty)$ and define $k\coloneqq 1+1/\delta$. Consider the following model for the Riesz representer:
$\alpha_{\bm{\beta}}(X)=\xi(X)\left(C+\left(1+\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\right)^{1/\delta}\right)-(1-\xi(X))\left(C+\left(1-\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\right)^{1/\delta}\right),$ (5)
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function and $\xi\colon{\mathcal{X}}\to\{0,1\}$ selects the branch. We call a link function in (5) a power link function.
@@
Under (5), the dual characterization implies that BP-Riesz regression returns the minimum BP-loss solution among approximately balancing models. In particular, if we use $\ell_{1}$-penalty in BP-Riesz regression, BP-Riesz regression is equivalent to solving a constrained problem of the form
@@
The power link (5) provides a continuous bridge between the specifications in the previous subsections. As $\delta\to 0$, we have $k=1+1/\delta\to\infty$ and
$\Big(1+\frac{t}{k}\Big)^{1/\delta}=\Big(1+\delta t+o(\delta)\Big)^{1/\delta}\to\exp(t),$
@@
so (5) reduces to the log link form used for UKL-Riesz regression. At $\delta=1$, the BP loss reduces to the squared loss, and the link becomes an affine transformation of the linear index around the origin, which connects BP-Riesz regression to SQ-Riesz regression up to reparameterization.
@@
This interpolation perspective is also consistent with the robustness interpretation of the BP divergence (Basu et al., 1998; Sugiyama et al., 2012). Smaller $\delta$ makes the objective closer to a KL-type criterion, which can be efficient under correct specification, while larger $\delta$ yields behavior closer to squared loss and is typically more robust to misspecification and extreme weights.
5 Automatic Neyman Orthogonalization
The advantage of (automatic) covariate balancing lies in the fact that it automatically implies Neyman orthogonality. Specifically, we consider an estimator given by
$\widehat{\theta}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\widehat{\alpha}(X_{i})Y_{i},$
where $\widehat{\alpha}$ is an estimator of the Riesz representer obtained from generalized Riesz regression. This estimator corresponds to the IPW estimator in ATE estimation (Horvitz & Thompson, 1952) and covariate shift adaptation with importance weighting (Shimodaira, 2000). We show that this estimator satisfies Neyman orthogonality if the regression function $\gamma_{0}$ belongs to the linear space spanned by ${\bm{\phi}}(X)$, which is used for Riesz representer estimation. We refer to this property as automatic Neyman orthogonalization.
5.1 Automatic Neyman Orthogonalization
The following result holds.
Theorem 5.1.
Consider generalized Riesz regression with the model
$\alpha(X)=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big),$
where $\zeta^{-1}$ is the inverse of a link function and ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a vector of basis functions. Suppose that $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$ holds. If $\gamma_{0}$ belongs to a linear space of ${\bm{\phi}}(X)$, then it holds that
$\widehat{\theta}$ $\coloneqq\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})Y_{i}$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-\gamma_{0}(X_{i})\big)+m(W_{i},\gamma_{0})\Big).$
If $\widehat{\alpha}$ satisfies the Donsker condition and is consistent, then $\widehat{\theta}$ is asymptotically efficient.
The derivation is as follows. Recall that the sample average of the Neyman orthogonal score is given by
$\frac{1}{n}\sum^{n}_{i=1}\psi(W;\widetilde{\eta},\theta)=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-\gamma_{0}(X_{i})\big)+m(W_{i},\gamma_{0})-\theta\Big),$
[…]
In Section 4, we model the Riesz representer as $\alpha(X)=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$. Then, a Riesz representer estimator $\widehat{\alpha}$ trained by generalized Riesz regression with $\lambda=0$ satisfies
$\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)=0$
for all $j=1,2,\dots,p$ by the automatic covariate balancing property (Corollary 4.2).
If $\gamma_{0}(X)$ is given as $\gamma_{0}(X)={\bm{\phi}}(X)^{\top}{\bm{\rho}}_{0}$ for some ${\bm{\rho}}_{0}\in{\mathbb{R}}^{p}$, then we have
$\frac{1}{n}\sum^{n}_{i=1}\psi(W;\widetilde{\eta},\theta)$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-\gamma_{0}(X_{i})\big)+m(W_{i},\gamma_{0})-\theta\Big),$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-{\bm{\rho}}_{0}^{\top}{\bm{\phi}}(X_{i})\big)+{\bm{\rho}}_{0}^{\top}m(W_{i},{\bm{\phi}}(X))-\theta\Big)$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})Y_{i}-\theta\Big),$
where we used
$\sum^{p}_{j=1}\rho_{0,j}\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)=0$
from the automatic covariate balancing property.
Such a property has been reported by Wong & Chan (2017) and Zhao (2019) in ATE estimation. Although they require a constant treatment effect, our results generalize them by allowing heterogeneous treatment effects. We emphasize that (i) the Riesz representer estimator $\widehat{\alpha}$ does not require a specific convergence rate; (ii) we cannot use cross fitting, and we assume the Donsker condition (if we use cross fitting, the covariate balancing collapses).
5.2 Automatic Neyman Error Minimization
Covariate balancing can also be interpreted as minimizing the estimation error of the Neyman orthogonal score. Given a nuisance parameter estimator $\widehat{\eta}$ and the true nuisance parameter $\eta_{0}$, the estimation error is given by
$\psi(W;\widehat{\eta},\theta)-\psi(W;\eta_{0},\theta)$
$=\Big(m(W,\widehat{\gamma})+\widehat{\alpha}(X)\big(Y-\widehat{\gamma}(X)\big)\Big)-\Big(m(W,\gamma_{0})+\alpha_{0}(X)\big(Y-\gamma_{0}(X)\big)\Big).$
Here, $\alpha_{0}(X)\big(Y-\gamma_{0}(X)\big)$ has mean zero and is ignorable in expectation. Therefore, we consider
$\text{NeymanError}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})\big(Y_{i}-\widehat{\gamma}(X_{i})\big)+m(W_{i},\widehat{\gamma})-m(W_{i},\gamma_{0})\right).$
Suppose that we estimate $\alpha_{0}$ by generalized Riesz regression using the basis function ${\bm{\phi}}(X)$, and that $\gamma_{0}$ belongs to the linear space spanned by ${\bm{\phi}}(X)$. Recall that we have
$\sum^{p}_{j=1}\rho_{0,j}\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)=0.$
If $\widehat{\gamma}$ equals $\gamma_{0}$ or $\widehat{\gamma}$ is estimated as ${\bm{\phi}}(X)^{\top}{\bm{\rho}}$, then by Neyman orthogonality, we have
$\text{NeymanError}=\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})Y_{i}-m(W_{i},\gamma_{0})\right).$
Furthermore, since $Y=\gamma_{0}(X)+\varepsilon$ for an error term $\varepsilon$ such that ${\mathbb{E}}[\varepsilon\mid X]=0$, we have
$\text{NeymanError}=\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})\gamma_{0}(X_{i})-m(W_{i},\gamma_{0})\right)+U,$
where $U$ is a term whose expectation is zero. Here, we again have
$\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})\gamma_{0}(X_{i})-m(W_{i},\gamma_{0})\right)=0$
from the automatic covariate balancing property. Thus, automatic covariate balancing and automatic Neyman orthogonalization also imply minimization of the Neyman error.
This property is discussed in Zhao (2019) for the case where UKL-Riesz regression with a basis function ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is used in ATE estimation. Our result generalizes the finding in that work to more general models, losses, and parameters of interest.
Cross fitting and the Neyman error
If we use cross fitting, the exact automatic covariate balancing property may not hold. On the other hand, exact automatic covariate balancing typically requires the Donsker condition to obtain an efficient estimator of the parameter of interest. We need to carefully address this trade-off.
Basis functions depending only on $Z$
In ATE estimation with standard propensity score modeling, we typically use basis functions depending on $Z$, that is, ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$. Under this choice, $m(W,\phi_{j}(\cdot))=0$. Therefore, automatic covariate balancing guarantees only $\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\alpha}(X_{i})\gamma_{0}(X_{i})\right)=0$. This result eliminates part of the Neyman error only when the conditional ATE given $x$ is invariant across $x$, that is, when the treatment effect is homogeneous. From the regression-function point of view, this requires that $\gamma_{0}(X)$ lies in the linear space spanned by ${\bm{\phi}}(Z)$. If we aim to guarantee automatic Neyman orthogonalization under treatment effect heterogeneity, we should use ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$, even though the Riesz representer corresponds to inverse propensity scores $1/e_{0}(Z)$ and $1/(1-e_{0}(Z))$.
5.3 Comparison with TMLE
TMLE is another promising approach in debiased machine learning (van der Laan, 2006). TMLE adds a perturbation to an initial estimate of $\gamma_{0}$. TMLE works to eliminate the sample average of the $(\star)$ part in the following Neyman error:
$\text{NeymanError}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\underbrace{\widehat{\alpha}(X_{i})(Y_{i}-\widehat{\gamma}(X_{i}))}_{\text{= $(\star)$}}+m(W_{i},\widehat{\gamma})-m(W_{i},\gamma_{0})\right).$
In contrast, generalized Riesz regression works to eliminate the sample average of the $(\star\star)$ part in the following Neyman error:
$\text{NeymanError}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})(Y_{i}-\underbrace{\widehat{\gamma}(X_{i}))+m(W_{i},\widehat{\gamma})}_{\text{= $(\star\star)$}}-m(W_{i},\gamma_{0})\right).$
That is, TMLE and generalized Riesz regression differ as follows:
  • TMLE puts the difficulty of efficient estimation of $\theta_{0}$ into the estimation of $\gamma_{0}$ by eliminating the influence of $\widehat{\alpha}$.
  • Generalized Riesz regression puts the difficulty of efficient estimation of $\theta_{0}$ into the estimation of $\alpha_{0}$ by eliminating the influence of $\widehat{\gamma}$.
Remark (TMLE).
We introduce TMLE in more detail. Let $\widehat{\gamma}^{(0)}$ be an initial estimate of $\gamma_{0}$. Then, given estimates $\widehat{\gamma}^{(0)}$ and $\widehat{\alpha}$, TMLE updates the regression function as
$\widehat{\gamma}^{(1)}(x)\coloneqq\widehat{\gamma}^{(0)}(x)+\frac{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})\big(Y_{i}-\widehat{\gamma}^{(0)}(X_{i})\big)}{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})^{2}}\widehat{\alpha}(x).$
It then estimates the parameter of interest as
$\widehat{\theta}^{\text{TMLE}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}m(W_{i},\widehat{\gamma}^{(1)}).$
Note that this update is derived as the solution in $\epsilon$ of
$\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\left(Y_{i}-\left(\widehat{\gamma}^{(0)}(X_{i})+\epsilon\widehat{\alpha}(X_{i})\right)\right)=0,$
which is given by
$\widehat{\epsilon}\coloneqq\frac{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})\{Y_{i}-\widehat{\gamma}^{(0)}(X_{i})\}}{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})^{2}}.$
Then, we derive $\widehat{\gamma}^{(1)}(x)$ as
$\widehat{\gamma}^{(1)}(x)=\widehat{\gamma}^{(0)}(x)+\widehat{\epsilon}\widehat{\alpha}(x).$
Under this update, the sample average of $(\star)$ in the Neyman error becomes zero.